A commercial jet can fly 550 mph on calm air. Traveling with the jet stream, the plane flew 2400 miles in the same amount of time it takes to fly 2000 miles against the jet stream. Find the rate of the jet stream.
Answer by checkley75(3666)

time=distance/rate. seeing as the times are the same we have the following equation:
2400/(550+r)=2000/(550-r) now cross multiply:
2400(550-r)=2000(550+r)
1,320,000-2400r=1,100,000+2000r
-2400r-2000r=1,100,000-1,320,000
-4400r=-220,000
r=-220,000/-4400
r=50 mph for the jet stream.
proof
2400/550+50)=2000/(550-50)
2400/600=2000/500
4=4